# Stability of low-momentum rest frame modes in first- and second-order spin hydrodynamics

### v1

@@ Abstract
The low-momentum rest frame stability of the second-order spin hydrodynamics is analyzed in detail. A truncation scheme of the Israel-Stewart formalism derived in our earlier work is proposed that extends the minimal causal formulation. It consists of altogether 40 interconnected relaxation-type dynamical equations – 16 (24) of them correspond to the independent components of the energy-momentum (spin) tensor. Similarly to previous studies, we find that the stability of the perturbations depends on the form of the spin equation of state. The stability is achieved if the equation of state satisfies the generalized Frenkel condition demanding that the “electric” and “magnetic” components of the spin density tensor have opposite signs. We then re-examine the study of the first-order (Navier-Stokes) dissipative spin hydrodynamics and find a similar type of behavior, which is consistent with the overall stability of low-momentum rest frame modes in conventional relativistic hydrodynamics.
@@ I Introduction
Investigations of the stability and causality properties of relativistic spin hydrodynamics have attracted a lot of attention lately [1, 2, 3, 4, 5, 6]. They shed light on the mathematical structure of the formalism and may determine its usefulness for description of realistic systems, for example, those produced in heavy-ion collisions [7, 8]. They can also be useful to distinguish between different formulations of spin hydrodynamics, the most popular of which are: the spin conserving approach [9, 10, 11, 12, 13, 14], the gradient expansion (of the first and second order [1, 15, 2, 16, 17, 18]) combined with the condition of positive entropy production, kinetic theory with non-local interactions [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], and quantum statistical mechanics [29, 30, 31, 32, 33, 34, 35], see also [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Stability analyses based on the gradient expansion use (in the leading order) the phenomenological form of the spin tensor [58]
@@
where $u^{\lambda}$ is the hydrodynamic flow and $S^{\mu\nu}$ is the spin density tensor. Moreover, the latter is taken in the form
@@
In this work we generalize the results obtained in Refs. [1, 2, 3, 4]. In Refs. [1, 2, 3] the first-order formulation was used, while in Ref. [4] the minimal causal spin hydrodynamics was studied. The latter is defined as an analog of the minimal causal extension of conventional hydrodynamics. It concentrates on the essential terms in the second order of gradient expansion to get a causal theory. Our present analysis can be treated as a generalization of the minimal causal spin hydrodynamics [4], as we consider the full 40 relaxation-type dynamical equations derived in [18] (with 16 (24) equations corresponding to the independent components of the energy-momentum (spin) tensor). The only truncation we propose is to neglect the so-called mixed terms. We systematically repeat the linear stability analysis of the low-momentum modes in the system’s rest frame and derive dispersion relations for all possible excitations. We find that the stability conditions for the extended model are essentially the same as those found before in the first-order and minimally causal theories. Hence, going to higher orders and including more terms in the second-order theory is not crucial for the stability of the low-momentum rest frame perturbations.
@@
Altogether, our analysis demonstrates that both the first- and second-order theories exhibit stability in the low-momentum rest frame provided the proper spin equation of state is used. This finding agrees with conventional relativistic hydrodynamics [59] and other relevant studies [60, 61, 62, 63, 64, 65, 66, 67]. Similar theoretical explorations employing linear mode analysis have also been conducted in the context of relativistic magneto-hydrodynamics [68, 69] and chiral hydrodynamics [70, 71]. While our focus has been on stability in the low-momentum regime, future studies could delve into the causality of this framework and its stability in the high-momentum limit. Furthermore, investigating stability at finite momentum values would be an intriguing endeavor, particularly in light of the stability concerns raised in a recent “minimal” study [4].
This work is organized as follows: Sec. II introduces our ”40=16+24” relaxation-time dynamical equations, laying the groundwork for the subsequent linear mode analysis presented in Sec. III. Our conclusions and outline of possible future studies are presented in Sec. IV.
@@ II Israel-Stewart-like Evolution Equation
III Stability Study
III.1 Linear mode analysis
@@
III.2 Coupled sound channels $\mathbf{A}$
@@
III.3 Degenerate coupled shear channels $\mathbf{B}$
@@
III.4 Coupled channels $\mathbf{C}$
@@
III.5 Decoupled non-propagating channels $\mathbf{D}$
@@
III.6 Stability Conditions

### v2

# Stability and causality of rest frame modes in second-order spin hydrodynamics

@@ Abstract
We analyze the low- and high-momentum rest frame modes in the second-order spin hydrodynamics and check the asymptotic causality of the theory. A truncation scheme of the Israel-Stewart formalism derived in our earlier work is proposed that extends the minimal causal formulation. It consists of altogether 40 interconnected relaxation-type dynamical equations – 16 (24) of them correspond to the independent components of the energy-momentum (spin) tensor. Similarly to previous studies, we find that the stability of the perturbations and asymptotic causality require using the spin equation of state that satisfies the generalized Frenkel condition demanding that the “electric” and “magnetic” components of the spin density tensor have opposite signs. For low-momentum modes this behavior is similar to that found earlier for the first-order (Navier-Stokes) spin hydrodynamics.
@@ I Introduction
Investigations of the stability and causality properties of relativistic spin hydrodynamics have attracted a lot of attention lately [1, 2, 3, 4, 5, 6]. They shed light on the mathematical structure of the formalism and may determine its usefulness for description of realistic systems, for example, those produced in heavy-ion collisions [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. They can also be useful to distinguish between different formulations of spin hydrodynamics, the most popular of which are: the spin conserving approach [19, 20, 21, 22, 23, 24], the gradient expansion (of the first and second order [1, 25, 2, 26, 27, 28]) combined with the condition of positive entropy production, kinetic theory with non-local interactions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and quantum statistical mechanics [39, 40, 41, 42, 43, 44, 45], see also [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Stability and causality analyses based on the gradient expansion use (in the leading order) the phenomenological form of the spin tensor [68]
@@
where $u^{\lambda}$ is the hydrodynamic flow and $S^{\mu\nu}$ is the spin density tensor. Recently, a connection between the phenomenological form (1) and the canonical formalism, where the spin tensor is totally antisymmetric, has been established [26]. Moreover, the spin density (1) is taken in the form
@@
In this work we generalize the results obtained in Refs. [1, 2, 3, 4]. In Refs. [1, 2, 3] the first-order formulation was used, while in Ref. [4] the minimal causal spin hydrodynamics was studied. The latter is defined as an analog of the minimal causal extension of conventional hydrodynamics. It concentrates on the essential terms in the second order of gradient expansion to get a causal theory. Our present analysis can be treated as a generalization of the minimal causal spin hydrodynamics, as we consider the full 40 relaxation-type dynamical equations derived in [28] (with 16 (24) equations corresponding to the independent components of the energy-momentum (spin) tensor). The only truncation we propose is to neglect the so-called mixed terms. We systematically repeat the linear stability analysis of the rest frame low- and high-momentum modes. Dispersion relations for all possible excitations are derived. We find that the stability conditions for the extended model are essentially the same as those found before in the first-order and minimally causal theories. Hence, going to higher orders and including more terms in the second-order theory is not crucial for the stability of the perturbations. Furthermore, we check the asymptotic causality of our framework and find that causality and stability conditions agree.
@@
Altogether, our analysis demonstrates that both the first- and second-order theories exhibit stability of the rest frame low-momentum modes provided the proper spin equation of state is used. This finding agrees with conventional relativistic hydrodynamics [69] and other relevant studies [70, 71, 72, 73, 74, 75, 76]. Moreover, the same conditions provide the high momentum stability and asymptotic causality of the considered framework (provided the used relaxation times are sufficiently large). Similar theoretical explorations employing linear mode analysis have also been conducted in the context of relativistic magneto-hydrodynamics [77] and chiral hydrodynamics [78, 79]. While we focus here on rest frame studies, our future work could delve into the stability, causality, and possible correlations between the two in a boosted frame. One novel approach leverages the information current introduced recently in Refs. [80, 81], especially considering its recent application in minimal spin hydrodynamics [82].
This work is organized as follows: Sec. II introduces our ”40=16+24” relaxation-time dynamical equations, which lays the foundation for the subsequent linear mode analysis in Sec. III. Sec. IV investigates the stability of the system in both low- and high-momentum regimes. Section V focuses on the asymptotic causality analysis. Finally, Sec. VI presents our conclusions and outlines potential areas for future research.
@@ II Israel-Stewart-like Evolution Equation
III Linear mode analysis
@@
III.1 Low-momentum Modes
III.1.1 Coupled sound channels $\mathbf{A}$
@@
III.1.2 Degenerate coupled shear channels $\mathbf{B}$
@@
III.1.3 Coupled channels $\mathbf{C}$
@@
III.1.4 Decoupled non-propagating channels $\mathbf{D}$
@@
III.2 High-momentum Modes
We now turn on to the characteristics of the modes within the high-momentum limit. It’s crucial to emphasize that the channel classification framework we have outlined in Section III.1 remains entirely valid in the high-momentum regime. From a physical standpoint, when the wavenumber along the z-direction, $k_{z}$, assumes large values, it becomes particularly insightful to prioritize our analysis on two categories of terms: those constituting the leading order and those exhibiting proportionality to the momentum itself. This allows us to extract the most essential information about the modes.
III.2.1 Coupled sound channels $\mathbf{A}$
Similar to the low-momentum case, the longitudinal fluctuations, vibrations parallel to the flow direction, are the focus of the sound channel. The sound channels can be identified with the fluctuations represented by the vector $\mathbf{v}_{\mathbf{A}}$ in Eq. (45). By analyzing the determinant of matrix $\mathbf{A}$ in Eq. (85) in the high-momentum limit we obtain the following modes,
${\omega_{1}=-\frac{i(\frac{4}{3}\eta+\xi)}{\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (108)
${\omega_{2}=-\frac{i(\frac{4}{3}\tilde{\chi}_{2}+\tilde{\chi}_{1})}{\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (109)
${\omega_{3,4}=-\frac{i}{2\tau_{\tau_{s}}\tau_{\Phi}}\frac{\tau_{\tau_{s}}^{2}\tilde{\chi}_{1}+\frac{4}{3}\tau_{\Phi}^{2}\tilde{\chi}_{2}}{\tau_{\tau_{s}}\tilde{\chi}_{1}+\frac{4}{3}\tau_{\Phi}\tilde{\chi}_{2}}\pm\sqrt{-\frac{\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}}{\tau_{\tau_{s}}\tau_{\Phi}}}\,k_{z}+\mathcal{O}(\frac{1}{k_{z}})},$ (110)
${\omega_{s_{1},s_{2}}=-i\frac{\sum\limits_{n=0}^{3}c_{n}v_{s_{1},s_{2}}^{2n}}{\sum\limits_{n=0}^{3}d_{n}v_{s_{1},s_{2}}^{2n}}+v_{s_{1},s_{2}}\,k_{z}+\mathcal{O}(\frac{1}{k_{z}})}.$ (111)
For the sake of simplicity, here, $\omega_{s_{1},s_{2}}$ represents 4 modes with $s_{1}=\pm$ and $s_{2}=\pm$. The 4 modes can be thought of as $\omega_{+,+},\,\omega_{+,-},\,\omega_{-,+},\,\omega_{-,-}$, where
${v_{s_{1},s_{2}}=s_{1}\sqrt{\frac{\mathcal{A}+s_{2}\mathcal{B}}{2(\varepsilon_{0}+p_{0})(\tau_{q}-\lambda^{\prime})\tau_{\pi}\tau_{\Pi}}}},$
${\mathcal{A}\equiv c_{s}^{2}(\varepsilon_{0}+p_{0})(\tau_{q}+3\lambda^{\prime})\tau_{\pi}\tau_{\Pi}+\tau_{q}(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi})},$
${\mathcal{B}^{2}\equiv\mathcal{A}^{2}-4c_{s}^{2}\tau_{\pi}\tau_{\Pi}\lambda(\tau_{q}-\lambda^{\prime})(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi})},$
along with the terms,
${c_{0}=-c_{s}^{2}\lambda^{\prime}\bigg{(}\tilde{\chi}_{1}\left(\frac{4\eta}{3}(\tau_{\Pi}+\tau_{\tau_{s}})+\xi(\tau_{\pi}+\tau_{\tau_{s}})\right)+\frac{4\tilde{\chi}_{2}}{3}\left(\frac{4\eta}{3}(\tau_{\Pi}+\tau_{\Phi})+\xi(\tau_{\pi}+\tau_{\Phi})\right)\bigg{)},}$
${c_{1}=\frac{1}{9}\bigg{(}-3c_{s}^{2}\bigg{(}\lambda^{\prime}\bigg{(}4\eta(\tau_{\Pi}\tau_{\Phi}+\tau_{\tau_{s}}(\tau_{\Pi}+\tau_{\Phi}))+3\xi(\tau_{\pi}\tau_{\Phi}+\tau_{\tau_{s}}(\tau_{\pi}+\tau_{\Phi}))}$
${\qquad-3(\varepsilon_{0}+p_{0})\left(\tau_{\pi}\tau_{\Pi}(3\tilde{\chi}_{1}+4\tilde{\chi}_{2})+4\tau_{\Phi}\tilde{\chi}_{2}(\tau_{\pi}+\tau_{\Pi})+3\tau_{\tau_{s}}\tilde{\chi}_{1}(\tau_{\pi}+\tau_{\Pi})\right)\bigg{)}}$
${-(\varepsilon_{0}+p_{0})\bigg{(}\tau_{q}\left(\tau_{\pi}\tau_{\Pi}(3\tilde{\chi}_{1}+4\tilde{\chi}_{2})+4\tau_{\Phi}\tilde{\chi}_{2}(\tau_{\pi}+\tau_{\Pi})+3\tau_{\tau_{s}}\tilde{\chi}_{1}(\tau_{\pi}+\tau_{\Pi})\right)+\tau_{\pi}\tau_{\Pi}(3\tau_{\tau_{s}}\tilde{\chi}_{1}+4\tau_{\Phi}\tilde{\chi}_{2})\bigg{)}\bigg{)}}$
${+4\eta\bigg{(}\tau_{q}(\tau_{\Pi}(3\tilde{\chi}_{1}+4\tilde{\chi}_{2})+3\tau_{\tau_{s}}\tilde{\chi}_{1}+4\tau_{\Phi}\tilde{\chi}_{2})+\tau_{\Pi}(3\tau_{\tau_{s}}\tilde{\chi}_{1}+4\tau_{\Phi}\tilde{\chi}_{2})\bigg{)}}$
${+3\xi\bigg{(}\tau_{q}(\tau_{\pi}(3\tilde{\chi}_{1}+4\tilde{\chi}_{2})+3\tau_{\tau_{s}}\tilde{\chi}_{1}+4\tau_{\Phi}\tilde{\chi}_{2})+\tau_{\pi}(3\tau_{\tau_{s}}\tilde{\chi}_{1}+4\tau_{\Phi}\tilde{\chi}_{2})\bigg{)}\bigg{)},}$
${c_{2}=\frac{1}{3}\bigg{(}(\varepsilon_{0}+p_{0})\bigg{(}3c_{s}^{2}\bigg{(}3\lambda^{\prime}(\tau_{\pi}\tau_{\Pi}\tau_{\Phi}+\tau_{\pi}\tau_{\tau_{s}}(\tau_{\Pi}+\tau_{\Phi})+\tau_{\Pi}\tau_{\tau_{s}}\tau_{\Phi})+\tau_{\pi}\tau_{\Pi}\tau_{q}(\tau_{\Phi}+\tau_{\tau_{s}})+\tau_{\tau_{s}}\tau_{\Phi}\left(\tau_{q}(\tau_{\pi}+\tau_{\Pi})+\tau_{\pi}\tau_{\Pi}\right)\bigg{)}}$
${+\lambda^{\prime}\bigg{(}\tau_{\pi}(3\tau_{\Pi}\tilde{\chi}_{1}+4\tau_{\Pi}\tilde{\chi}_{2}+4\tau_{\Phi}\tilde{\chi}_{2})+4\tau_{\Pi}\tau_{\Phi}\tilde{\chi}_{2}+3\tau_{\tau_{s}}\tilde{\chi}_{1}(\tau_{\pi}+\tau_{\Pi})\bigg{)}}$
${-\tau_{\pi}\tau_{\Pi}(4\tau_{\Phi}\tilde{\chi}_{2}+3\tau_{q}\tilde{\chi}_{1})-4\tau_{q}\tilde{\chi}_{2}\left(\tau_{\pi}(\tau_{\Pi}+\tau_{\Phi})+\tau_{\Pi}\tau_{\Phi}\right)-3\tau_{\tau_{s}}\tilde{\chi}_{1}\left(\tau_{q}(\tau_{\pi}+\tau_{\Pi})+\tau_{\pi}\tau_{\Pi}\right)\bigg{)}}$
${+4\eta\left(\tau_{\Pi}\tau_{q}\tau_{\Phi}+\tau_{q}\tau_{\tau_{s}}(\tau_{\Pi}+\tau_{\Phi})+\tau_{\Pi}\tau_{\tau_{s}}\tau_{\Phi}\right)+3\xi\left(\tau_{\pi}\tau_{q}\tau_{\Phi}+\tau_{q}\tau_{\tau_{s}}(\tau_{\pi}+\tau_{\Phi})+\tau_{\pi}\tau_{\tau_{s}}\tau_{\Phi}\right)\bigg{)},}$
${c_{3}=-(\varepsilon_{0}+p_{0})\bigg{(}-\lambda^{\prime}\left(\tau_{\pi}\tau_{\Pi}\tau_{\Phi}+\tau_{\tau_{s}}\tau_{\Phi}(\tau_{\pi}+\tau_{\Pi})+\tau_{\pi}\tau_{\Pi}\tau_{\tau_{s}}\right)+\tau_{\Phi}\left(\tau_{\Pi}\tau_{q}(\tau_{\pi}+\tau_{\tau_{s}})+\tau_{\pi}\tau_{q}\tau_{\tau_{s}}+\tau_{\pi}\tau_{\Pi}\tau_{\tau_{s}}\right)+\tau_{\pi}\tau_{\Pi}\tau_{q}\tau_{\tau_{s}}\bigg{)},}$
and,
${d_{0}=-2c_{s}^{2}\lambda^{\prime}\left(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi}\right)\left(\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}\right),}$
${d_{1}=4\bigg{(}\tau_{q}\left(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi}\right)\left(\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}\right)+c_{s}^{2}\left(-\lambda^{\prime}\tau_{\tau_{s}}\tau_{\Phi}\left(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi}\right)+(\varepsilon_{0}+p_{0})\tau_{\pi}\tau_{\Pi}(3\lambda^{\prime}+\tau_{q})\left(\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}\right)\right)\bigg{)},}$
${d_{2}=6\tau_{q}\tau_{\tau_{s}}\tau_{\Phi}\left(\frac{4}{3}\eta\tau_{\Pi}+\xi\tau_{\pi}\right)+6(\varepsilon_{0}+p_{0})\tau_{\pi}\tau_{\Pi}\left((\lambda^{\prime}-\tau_{q})\left(\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}\right)+c_{s}^{2}\tau_{\tau_{s}}\tau_{\Phi}(3\lambda^{\prime}+\tau_{q})\right),}$
${d_{3}=8(\varepsilon_{0}+p_{0})(\lambda^{\prime}-\tau_{q})\tau_{\tau_{s}}\tau_{\pi}\tau_{\Pi}\tau_{\Phi}.}$
III.2.2 Degenerate coupled shear channels $\mathbf{B}$
Shear channels deal with deviations perpendicular to the main flow, essentially orthogonal to the momentum direction $(0,0,k_{z})$. In similar process, yet in the high-momentum limit, as in (III.1) we obtain
${\omega_{1}=-\frac{i}{\tau_{q}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (112)
${\omega_{2}=-\frac{i(\gamma^{\prime}+\eta)}{\gamma^{\prime}\tau_{\pi}+\eta\tau_{\phi}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (113)
${\omega_{3}=-\frac{i(\tilde{\chi}_{2}+\tilde{\chi}_{3})}{\tau_{\tau_{a}}\tilde{\chi}_{2}+\tau_{\tau_{s}}\tilde{\chi}_{3}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (114)
${\omega_{4,5}=-\frac{i}{2}\left(\frac{1}{\tau_{\pi}}+\frac{1}{\tau_{\phi}}+\frac{\lambda^{\prime}}{\tau_{q}(\tau_{q}-\lambda^{\prime})}-\frac{\gamma^{\prime}+\eta}{\gamma^{\prime}\tau_{\pi}+\eta\tau_{\phi}}\right)\pm\sqrt{\frac{\tau_{q}\left(\gamma^{\prime}\tau_{\pi}+\eta\tau_{\phi}\right)}{(\varepsilon_{0}+p_{0})(\tau_{q}-\lambda^{\prime})\tau_{\pi}\tau_{\phi}}}\,k_{z}+\mathcal{O}(\frac{1}{k_{z}}),}$ (115)
${\omega_{6,7}=-\frac{i(\tau_{\tau_{a}}^{2}\tilde{\chi}_{2}+\tau_{\tau_{s}}^{2}\tilde{\chi}_{3})}{2\tau_{\tau_{a}}\tau_{\tau_{s}}(\tau_{\tau_{a}}\tilde{\chi}_{2}+\tau_{\tau_{s}}\tilde{\chi}_{3})}\pm\sqrt{-\frac{\tau_{\tau_{a}}\tilde{\chi}_{2}+\tau_{\tau_{s}}\tilde{\chi}_{3}}{\tau_{\tau_{a}}\tau_{\tau_{s}}}}\,k_{z}+\mathcal{O}(\frac{1}{k_{z}}),}$ (116)
${\omega_{8,9}=-\frac{i}{2\tau_{\Theta}}\pm\,\sqrt{\frac{\tilde{\chi}_{4}}{\tau_{\Theta}}}\,k_{z}+\mathcal{O}(\frac{1}{k_{z}}).}$ (117)
The shear channel modes are obtained from the condition $\textrm{det}(\mathbf{B})^{2}=0$, where the square implies that each shear mode in the group $\mathbf{B}$ is doubly degenerate.
III.2.3 Coupled channels $\mathbf{C}$
Channels $\mathbf{C}$ corresponding to new phenomena beyond relativistic hydrodynamics without spin. They are given by the vector $\mathbf{v}_{\mathbf{C}}$ defined in Eq. (45). Studying the determinant of the matrix $\mathbf{C}$ in Eq. (85), we obtain in the high-momentum momentum limit
${\omega_{1}=-\frac{i}{\tau_{\phi}}+\mathcal{O}(\frac{1}{k_{z}^{2}})+\mathcal{O}(\frac{1}{k_{z}^{4}}),}$ (118)
${\omega_{2,3}=-\frac{i}{2\tau_{\Theta}}\pm\sqrt{\frac{\tilde{\chi}_{4}}{\tau_{\Theta}}}k_{z}+\mathcal{O}(\frac{1}{k_{z}})+\mathcal{O}(\frac{1}{k_{z}^{2}}),}$ (119)
IV Stability Study